This issue is now confirmed resolved. Our team will be conducting maintenance to address the fault identified today.
Services should now be restored. If you are still experiencing issues, please attempt a restart of your router.
Our NOC team are continuing to monitor this situation closely.
Our team has identified the cause of the fault and remediation is in progress.
We are currently aware of transport issues to/from our 162G datacentre PoP. causing packet loss for some customers.
Our engineers are working on this as a top priority.